OpenStudy (anonymous):

Evaluate: [(40 + 5) − 32] ÷ 9 • 2 8.6 43 8 2

thomaster (thomaster):

Well the way i would solve this: \(\Large \dfrac{((40+5)-32)}{9}*2\) \(\Large \dfrac{45-32}{9}*2\) \(\Large \dfrac{13}{9}*2~=~\dfrac{26}{9}~≈~2.888\) But that's not in the options

OpenStudy (anonymous):

i know i cant figure it out

OpenStudy (anonymous):

it would be 13/18. which would be 0.7222

OpenStudy (anonymous):

i cant its not in the choices this is hard

thomaster (thomaster):

@amistre64 @ganeshie8 @Mertsj anyone of you knows how to solve this?

OpenStudy (amistre64):

40+5-32 then divide by 9 then multiply by 2

thomaster (thomaster):

Yea, so the answer is not in the options

OpenStudy (amistre64):

make sure you have the correct information posted

OpenStudy (amistre64):

is that 32 or 3^2?

OpenStudy (anonymous):

oh sorry guys its 3^2

OpenStudy (anonymous):

i feel so ad now

OpenStudy (amistre64):

:) s'ok

thomaster (thomaster):

Alright :) \(\Large \dfrac{45-3^2}{9}*2\) \(\Large \dfrac{36}{9}*2\)

OpenStudy (anonymous):

8

thomaster (thomaster):

Correct \(\huge \ddot\smile\)